OpenStudy (anonymous):

write as a single power of 3: 3 squareroot 3

OpenStudy (campbell_st):

\[3\sqrt{3} = \sqrt{3^2 \times 3} = \sqrt{27} = 27^(1/2)\]

OpenStudy (anonymous):

Sorry, That is right but its not what im looking for. Would it help if i mentioned it was for rational indices?

OpenStudy (campbell_st):

its 27^(1/2) or 27 to the power of 1/2

OpenStudy (campbell_st):

or how about 27 = (3^3) so square root 27 = (3^3) ^1/2 you should know how to do that